A taxi company had 157 riders in January of this year. In August they had 543 riders. Assuming the number of riders increased at a constant rate, write a linear equation to give the number of riders R in terms of the month M ( let the month of january be 1.) set rest of months values based on that starting pointt. Based on the equation how many riders did they have in April?
Answer by tutorcecilia(2152)

Let the months represent the x value and the ridership represent the y value.
January = x=1
February = x = 2
March = x = 3
April = 4, etc.
August = 8
.
So, the points are (1, 157) for January and (8, 543) for August.
Find the slope:
m = (543-157)/(8-1)=386/7 = 55.143
.
y=mx+b [Use the slope-intercept of a line]
157=(55.143)(1)+b [Plug-in the points and slope into the formula]
157=55.143+b [Solve for b]
101.86=b
.
y=mx+b [Plug-in the slope and y-intercept]
y=55.143x+101.86
.
To find the ridership in April, let x=4; plug the values into the slope-intercept formula and solve for y.